A fitness course claims that it can improve an individual's physical ability. To test the effect of a physical fitness course on one's physical ability, the number of sit-ups that a person could do in one minute, both before and after the course, was recorded. Ten individuals are randomly selected to participate in the course. The results are displayed in the following table. Can it be concluded, from the data, that participation in the physical fitness course resulted in significant improvement?

Let d=(number of sit-ups that can be done after taking the course)−(number of sit-ups that can be done prior to taking the course)d=(number of sit-ups that can be done after taking the course)−(number of sit-ups that can be done prior to taking the course). Use a significance level of α=0.05 for the test. Assume that the numbers of sit-ups are normally distributed for the population both before and after taking the fitness course.

Sit-ups before 22 47 25 23 41 45 24 49 28 41
Sit-ups after 33 55 34 27 47 57 29 53 45 48

Step 1 of 5: State the null and alternative hypotheses for the test.

Step 2 of 5:Find the value of the standard deviation of the paired differences. Round your answer to one decimal place.

Step 3 of 5:Compute the value of the test statistic. Round your answer to three decimal places

Step 4 of 5:Determine the decision rule for rejecting the null hypothesis H 0   . Round the numerical portion of your answer to three decimal places.

Step 5 of 5: Make the decision for the hypothesis test. (accept or reject null)
Answer #1

Given that,
null, H0: Ud = 0
alternate, H1: Ud != 0
level of significance, α = 0.05
from standard normal table, two tailed t α/2 =2.262
since our test is two-tailed
reject Ho, if to < -2.262 OR if to > 2.262
we use Test Statistic
to= d/ (S/√n)
where
value of S^2 = [ ∑ di^2 – ( ∑ di )^2 / n ] / ( n-1 ) )
d = ( Xi-Yi)/n) = -8.3
We have d = -8.3
pooled variance = calculate value of Sd= √S^2 = sqrt [ 841-(-83^2/10 ] / 9 = 4.111
to = d/ (S/√n) = -6.385
critical Value
the value of |t α| with n-1 = 9 d.f is 2.262
we got |t o| = 6.385 & |t α| =2.262
make Decision
hence Value of | to | > | t α| and here we reject Ho
p-value :two tailed ( double the one tail ) - Ha : ( p != -6.3846 ) = 0.0001
hence value of p0.05 > 0.0001,here we reject Ho
ANSWERS
---------------
null, H0: Ud = 0
alternate, H1: Ud != 0
test statistic: -6.385
critical value: reject Ho, if to < -2.262 OR if to > 2.262
decision: Reject Ho
p-value: 0.0001
we have enough evidence to support the claim that participation in the physical fitness course resulted in significant improvement